Grade 12Analytical Geometry

What is the area of region bounded by the curve y=x2 and line y=16 ?

Akshay chippigiri
8 Years agoGrade 12

1 Answer

Phanindra
8 Years ago
Y = x^2 and y =16....so.. X^2 = 16 so... x = +_ 4... So to find the area.... integrate y under the limits -4 to +4...As it is symmetric...Area = 2 × integral 0 to 4 x^2 dx . = 2 × x^3/3 limits are 0 to 4...Apply the limits to get the answer...So the answer is 128/3